Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
10x + 11y = -26.5
20x + 5y = -27.5

Answer by mananth(16949)

10 x + 11 y = -26.5 .............1
20 x + 5 y = -27.5 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 11
-50 x -55 y = 132.5
220 x 55 y = -302.5
Add the two equations
170 x = -170
/ 170
x = -1
plug value of x in (1)
10 x + 11 y = -26.5
-10 + 11 y = -26.5
11 y = -26.5 + 10
11 y = -16.5
y = -1.50